OpenStudy (anonymous):

25.568 is 37.6% of what number?

OpenStudy (anonymous):

\[25.568=0.376*x\] solve for x

OpenStudy (anonymous):

don't understand can you show me in steps

OpenStudy (anonymous):

25.568/x=37.6/100 Cross multiply and you get 2556.8=37.6x divide getting x=68

OpenStudy (anonymous):

thank u